How many pounds of candy worth $3.60/pound must be added to 10 pounds of chocolate worth $2.40/pound to form a mixture worth $3.30/pound?

Answer by Paul(988)

3.60x+2.40(10)=3.30(x+10)
3.60x+24=3.3x+33
0.6x=9
x=15

Hence, about 15 pounds should be added.
Paul.